Fermi National Laboratory
 

Accelerator Update for July 28, 2006 - July 31, 2006

 

Friday July 28

The day shift began with Operations monitoring store 4859 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, and with NuMI and MiniBooNE taking beam.

Operations began preparing the Recycler at 9:31 AM to stash antiprotons.

Operations turned off NuMI at 10:22 AM so that power supply technicians could repair the Lambertson filters (LAM60) and two power supplies (E:V108 and E:QF107).

Operations reported at 11:02 AM that high-voltage techs had repaired feeder 49 yesterday and switched power back to this feeder today.

At 1:52 PM, NuMI experts reported that they had found and begun repairs on the horn water skid and a horn power supply pressure switch.

FESS personnel reported at 3:17 PM that they were cleaning the Tevatron (TeV) sector A1 strainer. Earlier they had cleaned the A0 strainer.

At 3:58 PM, Operations attempted to turn NuMI on, but LAM61 tripped off on a filter rack summation indication. Power supply experts reported that there was no filter problem.

At 5:01 PM, Operations tried to turn NuMI on, but LAM61 again tripped off. A power supply expert disconnected the Lambertson from its filter.

Operations resumed sending beam to NuMI at 5:59 PM.

The Recycler began preparing to stash at 9:56 PM. The stash ended and Pbar resumed stacking at 10:34 PM.

Saturday July 29

The midnight shift began with Operations monitoring store 4859, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:04 AM, a Linac RF station (KRF4) tripped off. Operators had the station back up by 12:27 AM.

At 12:51 AM, Pbar tripped off due to a beam position problem. Operations fixed the problem by adjusting things in Booster.

At 2:01 AM, Operations reported that a beam control program (the Overthruster) was not working. An operator adjusted the timing for a transfer event and the program resumed normal operations.

At 3:26 AM, Operations turned off Pbar due to an Accumulator injection problem. The AP1 transfer line BPMs had failed. Operators rebooted the BPM system at 3:43 AM and reinstalled the parameter curves for a RF station (ARF1).

D0 experimenters reported at 4 AM that they were turning off their solenoid.

At 4:24 AM, Operations reported that the AP1 BPM readings were unrealistic and this was causing the Overthruster to move the beam all over. Operators bypassed one BPM (M:BPH104) and rebooted the rest. The ARF1 problem might have been due to the bad BPM readings. Later, an operator un-bypassed BPH104.

At 4:52 AM, the Linac Debuncher tripped off for the first of two times in about 39 minutes. It reset each time.

Operations terminated store 4859 at 5:35 AM.

Operations prepared DO and CDF for collision hall accesses at 5:52 AM.

At 10:06 AM, the NuMI beam permit tripped off due to a target and baffle flow alarm. Operators found nothing wrong. It reset okay.

At 12:03 PM, NuMI tripped off again on a target and baffle flow alarm. Operations informed an expert and then reset the beam permit.

At 2:19 PM, Booster suffered a trip due to beam position problems for the first of five times in one hour. Operators tuned most of the afternoon.

D0 came out of access at 3:33 PM.

CDF came out of access around 4:07 PM.

Pbar and MiniBooNE resumed taking beam at 4:39 PM.

At 4:40 PM, a NuMI kicker tripped off and would not reset. An operator called in an expert.

Operations put the TeV into a dry squeeze at 5:02 PM.

Operations put the TeV into shot setup at 6:01 PM.

At 6:17 PM, Operations put shot setup on hold due to a problem with a TeV sector D48 polarity switch. Operators investigated.

NuMI resumed taking beam at 6:22 PM. An expert put a fan on the kicker power supply.

At 6:22 PM, Operations reported that the quadrupole power supplies for all the low energy RF stations tripped off. Operators changed the master timing pulse module (TPM) and had Linac back up by 7:04 PM.

At 7:06 PM, Operations reported that Pbar stacked for a minute and then tripped off due to a pulsed magnet (PMag) interlocks trip.

Operators reported at 7:19 PM that they found the D48 polarity switch was down due to blown high voltage fuses. Operations called in a TeV expert.

At 7:22 PM, Operations reported that the temperature in the MI-60 kicker room was too high. Operators installed a fan, which lowered the room temperature to 97 F.

At 7:39 PM, Operations reported that LRF4 was producing many quadrupole power supply alarms. Operators installed a new TPM in LRF2 and then took the LRF2 TPM and installed it into LRF4.

At 8 PM, the cryo system techs at CHL reported potential water-cooling problem at D0. Operations contacted the Duty Mechanic.

Linac beam resumed at 8:13 PM.

At 8:33 PM, the power supply expert working on PMag reported that he looked at D:IB and that some of its filter capacitors had failed. Operations called in two more experts.

At 8:39 PM, a TeV expert and operators replaced the fuses for the D48 polarity switch, but they failed again. The expert repaired a loose connector and this fixed the problem.

Operations put the TeV into a dry squeeze at 8:58 PM.

Operations began the TeV shot setup at 9:28 PM.

At 11:18 PM, a Linac quadrupole power supply’s current (QPS105) was irregular. Operations called in a Linac expert.

Operations established store 4862 at 11:31 PM with an initial luminosity of 158.73E30.

At 11:44 PM, the Booster notcher tripped off. Operators contacted an expert.

Sunday July 30

The midnight shift began with Operations monitoring store 4862 and the stash, with Pbar, NuMI, and MiniBooNE off due to Linac and Booster problems.

Operators changed a control card for the Booster notcher at 12:32 AM.

The power supply expert reported that he replaced the resistor bank for D:IB at 12:52 AM and that it was back on line.

At 2:07 AM, the NuMI horn tripped off. It reset.

A Linac expert reported that he replaced the power supply for QPS105. Beam resumed at 3:40 AM.

The Recycler began preparing to stash at 4:08 AM. The stash ended at 4:35 AM.

At 7:29 AM, a power supply expert reported that he replaced a SCR pack for PMag, but there was still a problem in the tunnel. The magnet was found bad and has to be replaced. A crew was gathered to begin this process.

At 12:30 PM, the Duty Mechanic reported that he was concerned about with the Booster LCW temperatures. The Duty Electrician reported at 1:04 PM that he had found a chiller tripped off at CUB.

At 2:37 PM, an Accumulator power supply (A:IB) tripped off on a ground fault. The stack was lost. An operator contacted a power supply expert.

At 2:38 PM, Operations reported some Booster injection problems. An expert noted that the low level RF room temperature was high and had a high humidity. Operators installed a portable fan on the equipment racks and contacted the Duty Mechanic.

Operations reported at 4:19 PM that Booster temperatures were getting better.

At 6:17 PM, Operations reported that Pbar experts were searching for a ground fault on A:IB.

At 6:18 PM, L:QPS105 tripped off twice in ten minutes. Operators replaced a control card and tried to cool off the supply.

At 8:33 PM, Operations reported that a Recycler power supply (R:SQ704) had tripped off on a ground fault indication. Experts said they would look at this supply when they were finished with A:IB. The stash remained intact.

The power supply experts investigated R:SQ704, got advise from another expert, and then raised the ground fault detector voltage. The supply was back on by 9:09 PM.

At 9:41 PM, D0 experimenters reported that their solenoid was losing current. They had called in experts.

A Booster expert finished tuning the beam at 10:24 PM.

Pbar experts reported at 10:33 PM that they had completed the PMag replacement and that AP0 was ready for normal running.

D0 experts reported that they had found a problem with their scalers. The solenoid was back on by 11:05 PM.

Monday July 31

The midnight shift began with Operations monitoring store 4862 and the stash, with Pbar off due to A:IB work, and with NuMI and MiniBooNE taking beam.

At 3:29 AM, Pbar experts reported they had found what appeared to be a fault in the cables between the transformer and the supply or with the transformer itself. They went home to get some sleep.

At 4:09 AM, NuMI tripped off on a target and baffle flow alarm. Operations contacted an expert who said they could reset the alarm.

The Plans for Monday

The plans for today are to fix A:IB so that Pbar can resume stacking. Store 4862 will remain in the TeV until this evening. FESS reports that divers will continue to work on the A sector strainers. FESS also reports that they have invited a Marine Biologist to investigate the pond growth problems.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/31/2006    email Fermilab